OpenStudy (anonymous):

If an article is sold at 5% gain instead of 5% loss,the seller gets Rs 6.72 more. The C.P of the article is?

OpenStudy (anonymous):

So, s.p at 5%gain-s.p. at 5% loss=6.72

OpenStudy (anonymous):

(105/100*c.p.)-(95/100*c.p.)=6.72....solve for c.p.....

OpenStudy (anonymous):

c.p=67.2

OpenStudy (anonymous):

10/100c.p=6.72

OpenStudy (anonymous):

Yep....Nice work! U r correct!

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

cp=100/100+gain*s.p